# GaussianDWM: 3D Gaussian Driving World Model for Unified Scene Understanding and Multi-Modal Generation

### v1

@@ 1 Introduction
Driving World Model (DWM) [vista, gaia, drivingwm] have become essential for autonomous driving for their ability to predict future scene generation. These models predict environmental changes and synthesize simulation data for risk forecasting, route optimization, and corner case training. Most existing approaches achieve this by predicting modalities such as images [drivedreamer] and point clouds [pointcloud], which represent the visual and geometric properties of the environment. While these DWMs excel at forecasting how the environment may evolve, they are difficult to interpret, describe, or query, and cannot easily provide contextual information (e.g., visual question answering or scene description). With the rapid progress of Vision-Language Models (VLMs) [qwen3], remarkable advancements have been achieved in general vision tasks by leveraging world knowledge and causal reasoning. This highlights the potential of combining the scene understanding capability of VLMs with the generative power of DWMs as a promising future direction. A recent pioneering effort in this direction is HERMES [hermes]. It adopts BEV features to represent spatial information, aligns them with the text space, and incorporates them into the generative model. However, this BEV-based scene representation only achieves feature-level alignment between textual and spatial inputs, which is not sufficiently accurate. To overcome this limitation, we propose a novel 3DGS-based scene representation that unifies scene understanding and generation.
@@ 2 Related Work
Novel View Synthesis for Urban Scene With the emergence of NeRF [nerf] and 3D Gaussian Splatting (3DGS) [3dgs], many methods have adopted these representations across a wide range of tasks, including robotic mapping and localization [mneslam, mcnslam, plgslam, vpgsslam], VR [neslam, li2025human, gu2025mocount], and autonomous driving [nsg, snlidar]. Early NeRF-based methods such as NSG [nsg], SUDS [suds], ProSGNeRF [prosgnerf], EmerNeRF [emernerf], and FreeDriveRF [freedriverf] achieved dynamic–static disentangled reconstruction through the use of scene graphs, optical flow, or other motion cues. More recently, several 3DGS-based approaches have been proposed to further improve rendering efficiency. PVG [pvg] introduces periodic vibration-based temporal dynamics to unify static and dynamic elements without manual annotations. Methods such as Street Gaussians [streetgaussian], Driving Gaussian [drivinggaussian], and DeSiRe-GS [DeSiRe-GS] also explicitly separate dynamic and static components for reconstruction. LESSON [lesson] proposes a teacher-guided diffusion strategy for generating 3D Gaussian splats using only 2D supervision. STORM [storm] proposes a feed-forward Transformer architecture to infer dynamic Gaussians and their velocities, enabling efficient large-scale outdoor scene reconstruction. DrivingForward [drivingforward] achieves feed-forward reconstruction from sparse surround-view inputs using self-supervised pose and depth estimation. MUDG [mudg] and Dist-4D [dist4d] propose multi-modal novel view synthesis frameworks that generate both RGB and depth modalities.
@@
Driving World Model Driving world models [kong20253d, yan2025drivingsphere] have attracted considerable attention in autonomous driving due to their ability to provide comprehensive environmental representations and predict future scenarios. Current methods primarily rely on 2D and 3D conditions for scene generation. GAIA-1 [gaia1] introduces an autoregressive model for video generation in driving scenarios. DriveDreamer [drivedreamer] proposes a scene generation framework conditioned on 3D structure to provide geometric representations that benefit downstream autonomous driving tasks. MagicDrive [magicdrive] presents a street-view synthesis framework with precise 3D controls (e.g., camera poses, road maps) using cross-view attention, improving 3D object detection and BEV segmentation. DreamDrive [dreamdrive] further combines video diffusion with hybrid Gaussian scene representations to synthesize 4D scenes with 3D-consistent dynamic video rendering. Most recently, UniScene [uniscene] proposes an occupancy-centric approach that unifies semantic, visual, and LiDAR data generation. However, existing DWMs overlook the scene understanding ability of the driving environment.
@@
Large Language Models for Driving Large Language Models (LLMs) have demonstrated impressive generalization ability and extensive world knowledge across various tasks, including scene understanding [wang2025learning, 3dllm, 3dvla], visual question answering (VQA), and both 2D and 3D visual grounding. DriveGPT4 [drivegpt4] processes front-view video inputs to predict vehicle actions while providing natural language justifications via an LLM. DriveLM [drivelm] leverages LLMs for graph-based VQA and end-to-end autonomous driving. NuInteract [drivemonkey] further integrates large vision-language models (LVLMs) with a spatial processor using a set of learnable queries, trained on the large-scale NuInteract dataset containing over 1.5M multi-view image–language pairs, covering dense scene captioning and diverse interactive tasks. GaussianVLM [gaussianvlm] introduces a 3D Gaussian-based visual question answering (VQA) framework, where a SceneSplat-style variational autoencoder (VAE) is employed to directly encode 3D Gaussian scenes. Hermes [hermes] proposes a BEV-based world model that integrates scene representation with a Vision-Language Model (VLM) for joint scene understanding and generation, which is the most closely related work to ours. However, existing approaches typically rely on image, point cloud, or BEV representations. In contrast, our method leverages a 3D Gaussian scene representation [deng2025best3dscenerepresentation] that achieves explicit spatial alignment between language features and 3D geometry, resulting in more accurate multimodal correspondence and improved representation of both texture and structural information in complex environments.
@@ 3.1 World Tokenizer
3D Gaussian Tokenizer To preserve texture, 3D structure, and language alignment, we adopt 3D Gaussians as the scene representation for LLM input. We build upon LangSplat [langsplat] to construct a 3DGS language field, where each Gaussian is augmented with a language embedding $f_{i}$. These embeddings are obtained from CLIP features, which inherit hierarchical semantics extracted via SAM [sam]. We then follow the standard 3DGS rendering strategy, incorporating language information directly into the Gaussian primitives.
@@
3D Gaussian Projector We first align the extracted 3D Gaussian tokens to the text space. For each Gaussian primitive $G_{i}$, we represent its attributes as $G_{i}=(x_{i},o_{i},s_{i},r_{i},f_{i})$ where $x_{i}\in\mathbb{R}^{3}$ denotes the 3D spatial position, $o_{i}$ the opacity, $s_{i}$ the scale, $r_{i}$ the rotation, and $f_{i}$ the associated CLIP feature. For the Gaussian tokenizer, we first apply learnable Fourier embeddings [nerf] to the 3D coordinates $x_{i}$:
@@ 3.2 Scene Understanding
This section introduces the world understanding module. The Large Language Model (LLM) interprets driving scenarios from the world tokenizer outputs $\mathcal{G}_{i}\in\mathbb{R}^{N\times C}$ according to user instructions. Then, the LLM parses the user instruction $\mathcal{T}_{i}$ and extracts world knowledge from the driving scene, generating both a textual answer $t_{i}$ and a language feature representation $C_{L}$, which is later used as a condition signal for scene generation. This feature encodes high-level world knowledge as well as spatial information and is later used as a condition for scene generation. We implement the LLM using the widely adopted Qwen3 model [qwen3]. The overall architecture is:
@@ 4 Experiments
Table 1: The comparison between our GaussianDWM and other state-of-the-art models on the NuInteract dataset [drivemonkey]. The scene understanding task includes four subtasks: region description and perception, 2D visual grounding, 3D visual grounding, and planning. Our method achieves state-of-the-art results across all four tasks, which fully demonstrates the effectiveness of introducing a 3D Gaussian scene representation for enhancing the LLM’s capability to understand 3D spatial information.
@@ 4.1 Dataset and Evaluation Metric
We evaluate our method on two datasets for both scene understanding and scene generation. NuScenes [nuscenes] is a widely used autonomous driving dataset containing 700 training scenes, 150 validation scenes, and 150 test scenes. We use six surrounding camera images as inputs. NuInteract [drivemonkey] are recent benchmark for scene understanding. NuInteract provides $\sim$1.5M annotations and supports multiple tasks, including 2D perception and 3D visual grounding. For scene understanding, we adopt standard language and captioning metrics, including gCIDEr [cider], BLEU-4 [bleu], ROUGE [rouge]. For scene generation, we use FID for images, and FVD for videos as evaluation metrics.
@@ 4.2 Implementation Details
Our training pipeline consists of three stages. In the first stage, we train the Gaussian tokenizer, projector, and the proposed sampling strategy independently. We then integrate these components with the LLM and perform joint fine-tuning. For the LLM, we adopt a LoRA-based fine-tuning strategy. All models in this stage are trained using 16 NVIDIA A100 GPUs. In the second stage, we train the multi-modal generation module. We start by training a low-resolution ($224\times 400$) RGB video generation model, extend it to low-resolution RGB-D generation, and finally refine it into a high-resolution ($424\times 800$) RGB-D video generator using a mixed-frame-length strategy. The model is optimized using simulation-free rectified flow and a v-prediction loss [vprediction]. In the final stage, we perform end-to-end joint optimization over all components to ensure consistency between scene understanding and scene generation.
@@ 4.3 Scene Understanding
In this section, we present the results of scene understanding, which are divided into four tasks: region description and perception, planning, 2D visual grounding, and 3D visual grounding. The results are shown in Tab. 1. The experimental settings strictly follow Drivemonkey [drivemonkey]. We primarily compare two categories of methods: (1) LLM-based visual grounding approaches such as LLaVA [LLaVA] and InternVL [internvl], and (2) specialized 3D detection models such as BEVFormer [bevformer], PETR [petr], and CAPE [cape]. As shown in Tab. 1, GaussianDWM significantly outperforms previous LVLMs in terms of average metrics across all tasks on the NuInteract test dataset. It achieves state-of-the-art results on the RDP, 2D VG, and 3D VG tasks, effectively demonstrating the importance of introducing a 3D Gaussian scene representation for enabling LLMs to better comprehend 3D spatial information. Compared with previous point cloud– and BEV-based scene representations, our 3D Gaussian formulation allows for explicit alignment between 3D structural, texture, and language features, leading to more efficient and semantically consistent information injection into LLMs. Moreover, our task-aware hybrid sampling strategy efficiently selects Gaussian tokens most relevant to the query text while filtering out redundant 3D Gaussians. Compared with the current state-of-the-art VQA method Drivemonkey, our model shows clear advantages in both 2D and 3D visual grounding tasks, while also achieving comparable performance to specialized 3D detectors designed specifically for 3D VG.
@@ 4.4 Scene Generation
In this section, we present the evaluation of multi-modal scene generation, which includes both spatial scene generation and temporal scene generation. All experiments are conducted on the nuScenes dataset [nuscenes]. Following previous works [dist4d, magicdrive], we interpolate the 2Hz keyframe annotations to a higher frame rate of 12Hz, and the experimental settings strictly follow [dist4d].
Figure 4: Qualitative comparison of RGB-D NVS with 2m shift. Compared with state-of-the-art reconstruction-based methods for spatial NVS [omnire, deformable3dgs, pvg, streetgaussian], our method reduce artifacts of dynamic objects and preserves temporal-spatial consistency across large viewpoint shifts.
We compare our method with representative street-view synthesis approaches, including PVG [pvg], EmerNeRF [emernerf], StreetGaussian [streetgaussian], OmniRe [omnire], FreeVS [wang2025freevs], DiST-4D [Guo2025ICCV]. As shown in Tab. 3, our method outperforms all existing methods and achieves state-of-the-art performance. We use FID and FVD as evaluation metrics since ground-truth data is not available after spatial shifts. Our method demonstrates superior consistency and photorealism under extreme viewpoint shifts ($\pm 4$m), outperforming direct reconstruction-based approaches. This indicates that our framework effectively combines the advantages of 3D Gaussian scene representation and diffusion-based generative modeling. Under the guidance of our dual-condition mechanism—leveraging both high-level world knowledge and low-level geometric cues—our method achieves state-of-the-art spatial and texture fidelity under large viewpoint variations.
@@ Appendix B Dataset Details
To represent scenes from the NuScenes dataset [nuscenes] using 3D Gaussians, we construct the first large-scale 3D Gaussian dataset for outdoor urban environments, comprising 800 scenes derived from the original nuScenes sequences. Our goal is to build a dataset that enables generalizable 3D Gaussian–based scene understanding with LLMs, providing high-quality geometry and appearance representations suitable for downstream reasoning tasks.
Dataset Processing Many scenes contain frames with very low overlap or severe motion blur, making it difficult to obtain accurate 3DGS reconstructions. Therefore, we remove images or entire scenes whose reconstruction quality does not meet our standard. Since NuScenes dataset [nuscenes] provides high-quality LiDAR point clouds, we initialize each scene using aggregated LiDAR points, avoiding the large-scale inaccuracies that would arise from using COLMAP as in conventional 3DGS pipelines. For scenes with available depth supervision, we additionally apply a depth loss to improve geometric fidelity. After optimization, we further filter reconstructed scenes based on the PSNR metric before using them as inputs for our pre-training.
Language Label Collection and Training Our language label collection aims to establish 3D language paired data by associating each 3D Gaussian primitive with a rich language feature $f_{i}$. We build upon LangSplat [langsplat] to construct a 3DGS language field. These embeddings are obtained from CLIP features, which inherit hierarchical semantics extracted via SAM [sam]. We then follow the standard 3DGS rendering strategy, incorporating language information directly into the Gaussian primitives. As a collection of anisotropic 3D Gaussians, with each Gaussian $G(x)$ characterized by a mean $\mu\in\mathbb{R}^{3}$ and a covariance matrix $\Sigma$:

### v2

@@ 1 Introduction
Driving World Model (DWM) [17, 23, 48] have become essential for autonomous driving for their ability to predict future scene generation. These models predict environmental changes and synthesize simulation data for risk forecasting, route optimization, and corner case training. Most existing approaches achieve this by predicting modalities such as images [60] and point clouds [58], which represent the visual and geometric properties of the environment. While these DWMs excel at forecasting how the environment may evolve, they are difficult to interpret, describe, or query, and cannot easily provide contextual information (e.g., visual question answering or scene description). With the rapid progress of Vision-Language Models (VLMs) [55], remarkable advancements have been achieved in general vision tasks by leveraging world knowledge and causal reasoning. This highlights the potential of combining the scene understanding capability of VLMs with the generative power of DWMs as a promising future direction. A recent pioneering effort in this direction is HERMES [64]. It adopts BEV features to represent spatial information, aligns them with the text space, and incorporates them into the generative model. However, this BEV-based scene representation only achieves feature-level alignment between textual and spatial inputs, which is not sufficiently accurate. To overcome this limitation, we propose a novel 3DGS-based scene representation that unifies scene understanding and generation.
@@ 2 Related Work
Novel View Synthesis for Urban Scene With the emergence of NeRF [35] and 3D Gaussian Splatting (3DGS) [25], many methods have adopted these representations across a wide range of tasks, including robotic mapping and localization [12, 10, 11, 14], VR [13, 29, 18], and autonomous driving [36, 4]. Early NeRF-based methods such as NSG [36], SUDS [43], ProSGNeRF [8], EmerNeRF [57], and FreeDriveRF [49] achieved dynamic–static disentangled reconstruction through the use of scene graphs, optical flow, or other motion cues. More recently, several 3DGS-based approaches have been proposed to further improve rendering efficiency. PVG [5] introduces periodic vibration-based temporal dynamics to unify static and dynamic elements without manual annotations. Methods such as Street Gaussians [54], Driving Gaussian [63], and DeSiRe-GS [39] also explicitly separate dynamic and static components for reconstruction. LESSON [38] proposes a teacher-guided diffusion strategy for generating 3D Gaussian splats using only 2D supervision. STORM [56] proposes a feed-forward Transformer architecture to infer dynamic Gaussians and their velocities, enabling efficient large-scale outdoor scene reconstruction. DrivingForward [42] achieves feed-forward reconstruction from sparse surround-view inputs using self-supervised pose and depth estimation. MUDG [65] and Dist-4D [19] propose multi-modal novel view synthesis frameworks that generate both RGB and depth modalities.
@@
Driving World Model Driving world models [27, 53] have attracted considerable attention in autonomous driving due to their ability to provide comprehensive environmental representations and predict future scenarios. Current methods primarily rely on 2D and 3D conditions for scene generation. GAIA-1 [24] introduces an autoregressive model for video generation in driving scenarios. DriveDreamer [60] proposes a scene generation framework conditioned on 3D structure to provide geometric representations that benefit downstream autonomous driving tasks. MagicDrive [16] presents a street-view synthesis framework with precise 3D controls (e.g., camera poses, road maps) using cross-view attention, improving 3D object detection and BEV segmentation. DreamDrive [34] further combines video diffusion with hybrid Gaussian scene representations to synthesize 4D scenes with 3D-consistent dynamic video rendering. Most recently, UniScene [28] proposes an occupancy-centric approach that unifies semantic, visual, and LiDAR data generation. However, existing DWMs overlook the scene understanding ability of the driving environment.
@@
Large Language Models for Driving Large Language Models (LLMs) have demonstrated impressive generalization ability and extensive world knowledge across various tasks, including scene understanding [47, 22, 62], visual question answering (VQA), and both 2D and 3D visual grounding. DriveGPT4 [52] processes front-view video inputs to predict vehicle actions while providing natural language justifications via an LLM. DriveLM [41] leverages LLMs for graph-based VQA and end-to-end autonomous driving. NuInteract [61] further integrates large vision-language models (LVLMs) with a spatial processor using a set of learnable queries, trained on the large-scale NuInteract dataset containing over 1.5M multi-view image–language pairs, covering dense scene captioning and diverse interactive tasks. GaussianVLM [21] introduces a 3D Gaussian-based visual question answering (VQA) framework, where a SceneSplat-style variational autoencoder (VAE) is employed to directly encode 3D Gaussian scenes. Hermes [64] proposes a BEV-based world model that integrates scene representation with a Vision-Language Model (VLM) for joint scene understanding and generation, which is the most closely related work to ours. However, existing approaches typically rely on image, point cloud, or BEV representations. In contrast, our method leverages a 3D Gaussian scene representation [9] that achieves explicit spatial alignment between language features and 3D geometry, resulting in more accurate multimodal correspondence and improved representation of both texture and structural information in complex environments.
@@ 3.1 World Tokenizer
3D Gaussian Tokenizer To preserve texture, 3D structure, and language alignment, we adopt 3D Gaussians as the scene representation for LLM input. We build upon LangSplat [40] to construct a 3DGS language field, where each Gaussian is augmented with a language embedding $f_{i}$. These embeddings are obtained from CLIP features, which inherit hierarchical semantics extracted via SAM [26]. We then follow the standard 3DGS rendering strategy, incorporating language information directly into the Gaussian primitives.
@@
3D Gaussian Projector We first align the extracted 3D Gaussian tokens to the text space. For each Gaussian primitive $G_{i}$, we represent its attributes as $G_{i}=(x_{i},o_{i},s_{i},r_{i},f_{i})$ where $x_{i}\in\mathbb{R}^{3}$ denotes the 3D spatial position, $o_{i}$ the opacity, $s_{i}$ the scale, $r_{i}$ the rotation, and $f_{i}$ the associated CLIP feature. For the Gaussian tokenizer, we first apply learnable Fourier embeddings [35] to the 3D coordinates $x_{i}$:
@@ 3.2 Scene Understanding
This section introduces the world understanding module. The Large Language Model (LLM) interprets driving scenarios from the world tokenizer outputs $\mathcal{G}_{i}\in\mathbb{R}^{N\times C}$ according to user instructions. Then, the LLM parses the user instruction $\mathcal{T}_{i}$ and extracts world knowledge from the driving scene, generating both a textual answer $t_{i}$ and a language feature representation $C_{L}$, which is later used as a condition signal for scene generation. This feature encodes high-level world knowledge as well as spatial information and is later used as a condition for scene generation. We implement the LLM using the widely adopted Qwen3 model [55]. The overall architecture is:
@@ 4 Experiments
Table 1: The comparison between our GaussianDWM and other state-of-the-art models on the NuInteract dataset [61]. The scene understanding task includes four subtasks: region description and perception, 2D visual grounding, 3D visual grounding, and planning. Our method achieves state-of-the-art results across all four tasks, which fully demonstrates the effectiveness of introducing a 3D Gaussian scene representation for enhancing the LLM’s capability to understand 3D spatial information.
@@ 4.1 Dataset and Evaluation Metric
We evaluate our method on two datasets for both scene understanding and scene generation. NuScenes [3] is a widely used autonomous driving dataset containing 700 training scenes, 150 validation scenes, and 150 test scenes. We use six surrounding camera images as inputs. NuInteract [61] are recent benchmark for scene understanding. NuInteract provides $\sim$1.5M annotations and supports multiple tasks, including 2D perception and 3D visual grounding. For scene understanding, we adopt standard language and captioning metrics, including gCIDEr [44], BLEU-4 [37], ROUGE [31]. For scene generation, we use FID for images, and FVD for videos as evaluation metrics.
@@ 4.2 Implementation Details
Our training pipeline consists of three stages. In the first stage, we train the Gaussian tokenizer, projector, and the proposed sampling strategy independently. We then integrate these components with the LLM and perform joint fine-tuning. For the LLM, we adopt a LoRA-based fine-tuning strategy. All models in this stage are trained using 16 NVIDIA A100 GPUs. In the second stage, we train the multi-modal generation module. We start by training a low-resolution ($224\times 400$) RGB video generation model, extend it to low-resolution RGB-D generation, and finally refine it into a high-resolution ($424\times 800$) RGB-D video generator using a mixed-frame-length strategy. The model is optimized using simulation-free rectified flow and a v-prediction loss [15]. In the final stage, we perform end-to-end joint optimization over all components to ensure consistency between scene understanding and scene generation.
@@ 4.3 Scene Understanding
In this section, we present the results of scene understanding, which are divided into four tasks: region description and perception, planning, 2D visual grounding, and 3D visual grounding. The results are shown in Tab. 1. The experimental settings strictly follow Drivemonkey [61]. We primarily compare two categories of methods: (1) LLM-based visual grounding approaches such as LLaVA [32] and InternVL [6], and (2) specialized 3D detection models such as BEVFormer [30], PETR [33], and CAPE [51]. As shown in Tab. 1, GaussianDWM significantly outperforms previous LVLMs in terms of average metrics across all tasks on the NuInteract test dataset. It achieves state-of-the-art results on the RDP, 2D VG, and 3D VG tasks, effectively demonstrating the importance of introducing a 3D Gaussian scene representation for enabling LLMs to better comprehend 3D spatial information. Compared with previous point cloud– and BEV-based scene representations, our 3D Gaussian formulation allows for explicit alignment between 3D structural, texture, and language features, leading to more efficient and semantically consistent information injection into LLMs. Moreover, our task-aware hybrid sampling strategy efficiently selects Gaussian tokens most relevant to the query text while filtering out redundant 3D Gaussians. Compared with the current state-of-the-art VQA method Drivemonkey, our model shows clear advantages in both 2D and 3D visual grounding tasks, while also achieving comparable performance to specialized 3D detectors designed specifically for 3D VG.
@@ 4.4 Scene Generation
In this section, we present the evaluation of multi-modal scene generation, which includes both spatial scene generation and temporal scene generation. All experiments are conducted on the nuScenes dataset [3]. Following previous works [19, 16], we interpolate the 2Hz keyframe annotations to a higher frame rate of 12Hz, and the experimental settings strictly follow [19].
Figure 4: Qualitative comparison of RGB-D NVS with 2m shift. Compared with state-of-the-art reconstruction-based methods for spatial NVS [7, 59, 5, 54], our method reduce artifacts of dynamic objects and preserves temporal-spatial consistency across large viewpoint shifts.
We compare our method with representative street-view synthesis approaches, including PVG [5], EmerNeRF [57], StreetGaussian [54], OmniRe [7], FreeVS [45], DiST-4D [20]. As shown in Tab. 3, our method outperforms all existing methods and achieves state-of-the-art performance. We use FID and FVD as evaluation metrics since ground-truth data is not available after spatial shifts. Our method demonstrates superior consistency and photorealism under extreme viewpoint shifts ($\pm 4$m), outperforming direct reconstruction-based approaches. This indicates that our framework effectively combines the advantages of 3D Gaussian scene representation and diffusion-based generative modeling. Under the guidance of our dual-condition mechanism—leveraging both high-level world knowledge and low-level geometric cues—our method achieves state-of-the-art spatial and texture fidelity under large viewpoint variations.
@@ 5 Conclusion
Acknowledgement We sincerely thank Yikang Ding from Kuaishou Kling AI for his kind help and support.
@@ Appendix B Dataset Details
To represent scenes from the NuScenes dataset [3] using 3D Gaussians, we construct the first large-scale 3D Gaussian dataset for outdoor urban environments, comprising 800 scenes derived from the original nuScenes sequences. Our goal is to build a dataset that enables generalizable 3D Gaussian–based scene understanding with LLMs, providing high-quality geometry and appearance representations suitable for downstream reasoning tasks.
Dataset Processing Many scenes contain frames with very low overlap or severe motion blur, making it difficult to obtain accurate 3DGS reconstructions. Therefore, we remove images or entire scenes whose reconstruction quality does not meet our standard. Since NuScenes dataset [3] provides high-quality LiDAR point clouds, we initialize each scene using aggregated LiDAR points, avoiding the large-scale inaccuracies that would arise from using COLMAP as in conventional 3DGS pipelines. For scenes with available depth supervision, we additionally apply a depth loss to improve geometric fidelity. After optimization, we further filter reconstructed scenes based on the PSNR metric before using them as inputs for our pre-training.
Language Label Collection and Training Our language label collection aims to establish 3D language paired data by associating each 3D Gaussian primitive with a rich language feature $f_{i}$. We build upon LangSplat [40] to construct a 3DGS language field. These embeddings are obtained from CLIP features, which inherit hierarchical semantics extracted via SAM [26]. We then follow the standard 3DGS rendering strategy, incorporating language information directly into the Gaussian primitives. As a collection of anisotropic 3D Gaussians, with each Gaussian $G(x)$ characterized by a mean $\mu\in\mathbb{R}^{3}$ and a covariance matrix $\Sigma$: